OpenStudy (anonymous):

What is cos (x/2) =0?

OpenStudy (anonymous):

You are trying to find when cos(x/2) is zero. What angles have a cosine of zero? \[\frac{ \pi }{ 2 } and \frac{ 3\pi }{ 2 } \] will give you a cosine of zero. So, set these two angles equal to x/2 and solve for x. \[\frac{ x }{ 2 }=\frac{ \pi }{ 2 } \] so \[x=\pi\] Similarly, \[\frac{ x }{ 2 }=\frac{ 3\pi }{ 2 }\] so \[x=3\pi \]